# Marginal treatment effects in the absence of instrumental variables

### v2

@@ 3 Identification
Assumption CMI (Conditional Mean Independence). Assume that $E\left[U_{1}-U_{0}\left|V,X\right.\right]=E\left[U_{1}-U_{0}\left|V\right.\right]$ with probability one.
Assumption CMI is standard in the MTE literature and commonly referred to as separability or additive separability assumption (e.g., Brinch et al., 2017; Mogstad et al., 2018; Zhou and Xie, 2019), because it’s a necessary and sufficient condition for the MTE to be additively separable into observables and unobservables (Zhou and Xie, 2019, p.3074):
$\Delta^{\text{MTE}}\left(x,v\right)$ $=$ $h_{1}\left(x\right)-h_{0}\left(x\right)+E\left[U_{1}-U_{0}\left|X=x,V=v\right.\right]$
$=$ $h_{1}\left(x\right)-h_{0}\left(x\right)+E\left[U_{1}-U_{0}\left|V=v\right.\right].$
@@
where $\mu_{U\left|X\right.}\left(x\right)=\left(\sigma_{UX}\left/\sigma_{X}^{2}\right.\right)x$ and $\sigma_{U\left|X\right.}^{2}=1-\left(\sigma_{UX}^{2}\left/\sigma_{X}^{2}\right.\right)$. Hence,
@@
$E\left[Y\left|X,D=d\right.\right]$ $=$ $E\left[Y_{d}\left|X,D=d\right.\right]$ (5)
$=$ $h_{d}\left(X\right)+E\left[U_{d}\left|X,D=d\right.\right]$
@@
[…]
By multiplying both sides of the above equations with $p$ or $1-p$ in order to eliminate the denominators and then differentiating with respect to $p$, we obtain
$g_{1}\left(p\right)+pg_{1}^{\left(1\right)}\left(p\right)$ $=$ $E\left[\left.U_{1}\right|V=p\right],$
$-g_{0}\left(p\right)+\left(1-p\right)g_{0}^{\left(1\right)}\left(p\right)$ $=$ $-E\left[\left.U_{0}\right|V=p\right],$
@@
Recall that $P=\pi\left(X\right)$, where $\pi$ is a deterministic function, albeit unknown. Given that the argument of $g_{d}$ in (5) exhibits no variation apart from that generated by $X$, it is impossible to distinguish between $g_{d}$ and $h_{d}$ in the absence of additional assumptions. For example, one may choose to let $h_{d}$ absorb $g_{d}\left(\pi\left(\cdot\right)\right)$ and set $g_{d}=0$, or vice versa. Functional form restrictions that differentiate $h_{d}$ and $g_{d}\left(\pi\left(\cdot\right)\right)$ from one another can address this issue and facilitate the identification of both $h_{d}$ and $g_{d}$. To this end, we impose linearity on $h_{d}$ and nonlinearity on $\pi$.
@@
Assumption L imposes a linear restriction on the potential outcome regression functions, which is a common practice in empirical studies. The linear restriction may seem too strong compared with that in the existing identification strategies, which allow highly flexible (especially nonparametric) specifications on the potential outcomes. However, for the estimation of MTE, the linear specification has been nearly universally adopted for tractability and interpretability (e.g., Kirkeboen et al., 2016; Kline and Walters, 2016; Brinch et al., 2017; Heckman et al., 2018; Mogstad et al., 2021; Aryal et al., 2022; Mountjoy, 2022). Noting that Assumption L implicitly requires the potential outcomes to be continuously distributed and supported on the entire real line, we would like to point out that our identification strategy can also be adapted to the case of limited valued outcomes by specifying linear latent index. A detailed discussion is left to Appendix C. With a little abuse of notation, we redenote $U_{d}$ to absorb the intercept $\alpha_{d}$ so that
@@
$E\left[Y\left|X,D=d\right.\right]$ $=$ $X^{\prime}\beta_{d}+g_{d}\left(P\right),\text{ }d=0,1,$ (7)
@@
Assumption NL (Non-Linearity). Assume that $\pi_{0}\left(x^{C}\right)$ and $E\left[Y\left|X^{C}=x^{C},X^{D}=0,D=d\right.\right]$, $d=0,1$, are differentiable with respect to $x^{C}$, and that the derivative of $\pi_{0}$ satisfies the following NL2 when $\dim\left(X^{C}\right)\geq 2$ or NL1 when $\dim\left(X^{C}\right)=1.$
@@
Assumption NL requires the propensity score function to be nonlinear in continuous covariates in a generalized sense. The combination of Assumptions L and NL will enable identification based on functional form in a semiparametric version, which can realize the identification of linear coefficients by exploiting nonlinearity elsewhere in the model. The nonlinearity assumption has two different forms, depending on the number of continuous covariates. When two or more continuous covariates are available, Assumption NL2 will require some variation in $\pi_{0}$ to distinguish it from a linear-index function. Concretely, NL2 will not hold if $\pi_{0}\left(x^{C}\right)=f\left(x^{C\prime}\gamma\right)$ for a smooth function $f$, because in this case, both sides of the inequality in (v) are equal to $\left.\gamma_{k}\right/\gamma_{j}$. Otherwise, however, it is difficult to construct examples that violate (v). In practice, (v) can be fulfilled even when $\pi_{0}$ is single-index specified, if interaction or/and quadratic terms are added, as shown in Example 4.
@@
Assumption NL2 requires the existence of at least two continuous covariates. However, in empirical studies based on survey data, most of the demographic characteristics are documented as discrete or categorical variables, such as age, gender, race, marital status, educational attainment, and so on. Therefore, we also impose Assumption NL1, as a supplement to NL2, to take into account the situation in which only one continuous covariate is available. NL1 requires $\pi_{0}$ to have a stationary point, which implies that $\pi_{0}$ is necessarily nonlinear. NL1 will hold if the probability of receiving treatment is unaffected by some local change in the continuous covariate. It’s a differential version of the local irrelevance assumption imposed in nonseparable models for attaining point identification (e.g., Torgovitsky, 2015; D’Haultfœuille et al., 2021). NL1 can be extended to the case of no continuous covariate, as discussed in Appendix B. However, using only discrete covariates provides little identifying variation, which may lead to poor performance in the subsequent model estimation (Garlick and Hyman, 2022). Hence, we focus on the case of at least one continuous covariate.
@@
$D=1\left\{\frac{X^{\prime}\gamma}{\sigma\left(X\right)}\geq\tilde{U}\right\},$
that is, the reduced-form error $V$ in (3) is equal to $\tilde{U}$, and the propensity score function is equal to $\pi\left(x\right)=\left.\left(x^{\prime}\gamma\right)\right/\sigma\left(x\right)$. In general, $\pi\left(x\right)$ is a nonlinear function. Another specification we consider is a linear-index model with endogeneity in a certain component $X^{e}$ of $X^{C}$, in which $\pi\left(x\right)=F_{U\left|X^{e}\right.}\left(\left.x^{\prime}\gamma\right|x^{e}\right)$. Given that $\pi\left(x\right)$ is generally highly nonlinear in $x^{e}$, Assumption NL1 holds straightforwardly, and NL2 holds with $x_{k}^{C}=x^{e}$.
Unlike in the binary response model, the ubiquitous heteroscedasticity and endogeneity benefit our results while inducing no trouble, because the identification of the MTE is irrelevant to the structural coefficients in $\gamma$. Our MTE is defined by the reduced-form treatment error; thus, all we need from the treatment equation is the propensity score, which has a reduced-form nature.
@@
The proof of Theorem 1 is grounded on the observed regression functions (7), which summarize the information from the data. As $g_{d}$ is unknown, we need to eliminate it through some subtraction to realize the identification of $\beta_{d}$. When only one continuous covariate exists, the subtraction can be carried out locally around the point satisfying Assumption NL1. Otherwise, our strategy is to perturb two continuous covariates, such as $x_{k}^{C}$ and $x_{j}^{C}$, in such a way that $\pi\left(x\right)$ remains unchanged. Specifically, for each group $d$, we increase $x_{k}^{C}$ by a small $\epsilon$ and simultaneously change $x_{j}^{C}$ by $\epsilon$ multiplied by $\left.-\partial_{k}\pi_{0}\left(x^{C}\right)\right/\partial_{j}\pi_{0}\left(x^{C}\right)$, resulting in a perturbed value of the regression function. Subtracting the perturbed regression function from the original (7) will cancel out $g_{d}$ due to the equality of $\pi\left(x\right)$, giving rise to an equation for $\beta_{d,k}^{C}$ and $\beta_{d,j}^{C}$. Note that the multiplier $\left.-\partial_{k}\pi_{0}\left(x^{C}\right)\right/\partial_{j}\pi_{0}\left(x^{C}\right)$ is the partial derivative of $x_{j}^{C}$ with respect to $x_{k}^{C}$ if we view $x_{j}^{C}$ as an implicit function of the other continuous covariates by equating $\pi_{0}\left(x^{C}\right)$ to a constant. Accordingly, Assumption NL2.(v) implies two linearly independent equations, ensuring an exact solution (i.e., identification) of $\beta_{d,k}^{C}$ and $\beta_{d,j}^{C}$. The detailed proof of Theorem 1 is presented in Appendix A. It is worth mentioning that our strategy naturally features overidentification in the sense that generally more than one pair of points in the support of $X^{C}$ satisfy Assumption NL2 or more than one point satisfies NL1, because $X^{C}$ is continuously distributed. Moreover, there may be more than one value of $X^{D}$ that satisfies Assumptions NL and S. Consequently, the identification can also be represented as the average of solutions over all points of $X^{C}$ satisfying Assumption NL and over all values of $X^{D}$ satisfying Assumptions NL and S.
@@ 4 Estimation
Our identification strategy for the MTE implies a separate estimation procedure that works with the partially linear regression (7) by each treatment status. In particular, we recommend the kernel-weighted pairwise difference estimation method proposed by Ahn and Powell (1993) because of its computational simplicity, well-established asymptotic properties, and, most importantly, its capacity to effectively leverage the overidentifying information that underlies the data. Suppose that $\left\{\left(Y_{i},D_{i},X_{i}\right):i=1,2,\cdots,n\right\}$ is a random sample of observations on $\left(Y,D,X\right)$. In the first step, we estimate the nonparametrically specified propensity score using the kernel method, that is,
$\hat{\pi}\left(x\right)=\frac{\sum_{i=1}^{n}D_{i}\left[\prod_{l=1}^{\dim\left(X^{C}\right)}k_{1}\left(\left.\left(X_{il}^{C}-x_{l}^{C}\right)\right/h_{1l}\right)\right]1\left\{X_{i}^{D}=x^{D}\right\}}{\sum_{i=1}^{n}\left[\prod_{l=1}^{\dim\left(X^{C}\right)}k_{1}\left(\left.\left(X_{il}^{C}-x_{l}^{C}\right)\right/h_{1l}\right)\right]1\left\{X_{i}^{D}=x^{D}\right\}}$ (10)
and
$\hat{P}_{i}=\hat{\pi}\left(X_{i}\right),$ (11)
where $h_{1l}$, $l=1,2,\cdots,\dim\left(X^{C}\right)$, are bandwidths and $k_{1}$ is a univariate kernel function. If the dimension of $X^{D}$ is large, then a smoothed kernel for discrete covariates (Racine and Li, 2004) can be applied as a substitute for the indicator function, to alleviate the potential problem of inadequate observations in each data cell divided by the support of $X^{D}$. If the number of continuous covariates is not small either, then the well-known curse of dimensionality will appear, and a linear-index specification may thus be practically more relevant when modeling the propensity score. The index should include a series of interaction terms and quadratic or even higher-order terms of continuous covariates to supply sufficient nonlinear variation. The linear-index propensity score can be estimated by parametric probit/logit or semiparametric methods (e.g., Powell et al., 1989; Ichimura, 1993; Klein and Spady, 1993; Lewbel, 2000), depending on the distributional assumption on the error term.
@@
$\hat{\omega}_{dij}=1\left\{D_{i}=D_{j}=d\right\}k_{2}\left(\frac{\hat{P}_{i}-\hat{P}_{j}}{h_{2}}\right),$
@@
$\hat{w}_{di}\left(p\right)=1\left\{D_{i}=d\right\}k_{3}\left(\frac{\hat{P}_{i}-p}{h_{3}}\right).$
Finally, we plug $\hat{\beta}_{d}$, $\hat{g}_{d}$, $\hat{g}_{d}^{\left(1\right)}$, and $\hat{\pi}$ into identification equations (8) and (9) to estimate the MTE and other causal parameters, as follows:
$\hat{\Delta}^{\text{MTE}}\left(x,v\right)$ $=$ $x^{\prime}\left(\hat{\beta}_{1}-\hat{\beta}_{0}\right)+\left[\hat{g}_{1}\left(v\right)-\hat{g}_{0}\left(v\right)\right]+v\hat{g}_{1}^{\left(1\right)}\left(v\right)+\left(1-v\right)\hat{g}_{0}^{\left(1\right)}\left(v\right),$

### v3

@@ 3 Identification
[…]
Assumption CMI is standard in the MTE literature and commonly referred to as separability or additive separability assumption (e.g., Brinch et al., 2017; Mogstad et al., 2018; Zhou and Xie, 2019), because it’s a necessary and sufficient condition for the MTE to be additively separable in observables and unobservables (Zhou and Xie, 2019, p.3074):
[…]
@@
[…]
@@
[…]
@@
$g_{1}\left(p\right)$ $=$ $E\left[\left.U_{1}\right|V\leq p\right]=\frac{1}{p}\int_{0}^{p}E\left[\left.U_{1}\right|V=v\right]dv,$
$g_{0}\left(p\right)$ $=$ $E\left[\left.U_{0}\right|V>p\right]=\frac{1}{1-p}\int_{p}^{1}E\left[\left.U_{0}\right|V=v\right]dv.$
The function $g_{d}\left(\cdot\right)$ is usually referred to as selectivity bias correction term and plays a key part in the literature on sample selection models. We introduce $g_{d}\left(\cdot\right)$ into the identification of MTE because $g_{d}\left(\cdot\right)$ is closely related to the MTE and it is generally easier to identify $g_{d}\left(\cdot\right)$ than directly identify the MTE. From its definition, $g_{d}\left(\cdot\right)$ can be viewed as an aggregating function of the unobservable heterogeneity of the MTE. By multiplying $g_{d}\left(p\right)$ with $p$ or 1p1-p and then differentiating with respect to $p$, we can retrieve the unobservable part of the MTE as
$E\left[\left.U_{1}\right|V=p\right]$ $=$ $g_{1}\left(p\right)+pg_{1}^{\left(1\right)}\left(p\right),$
$E\left[\left.U_{0}\right|V=p\right]$ $=$ $g_{0}\left(p\right)-\left(1-p\right)g_{0}^{\left(1\right)}\left(p\right),$
@@
Recall that $P=\pi\left(X\right)$, where $\pi$ is a deterministic function, albeit unknown. Without having an IV, especially without having the exclusion restriction that assumes some element of $X$ to drop out of $h_{d}\left(X\right)$, the argument of $g_{d}$ in (6) exhibits no extra variation than $h_{d}$. Therefore, it is impossible to distinguish between $g_{d}$ and $h_{d}$. For example, one may choose to let $h_{d}\left(\cdot\right)$ absorb $g_{d}\left(\pi\left(\cdot\right)\right)$ and set $g_{d}=0$. Inspired by the identification result of Escanciano et al. (2016) for a double index model, functional form restrictions that differentiate $h_{d}\left(\cdot\right)$ and $g_{d}\left(\pi\left(\cdot\right)\right)$ from one another can address this issue and facilitate the identification of both $h_{d}$ and $g_{d}$. To this end, we impose linearity on $h_{d}$ and nonlinearity on $\pi$.
@@
Assumption L imposes a linear restriction on the potential outcome regression functions, which is a common practice in empirical studies. The linear restriction may seem too strong compared with the existing identification strategies for the MTE, which generally allow highly flexible (especially nonparametric) specifications on the potential outcomes. However, when the identification is accomplished and the estimation comes into consideration, the linear specification will turn to be nearly universally adopted due to its tractability and interpretability (e.g., Kirkeboen et al., 2016; Kline and Walters, 2016; Brinch et al., 2017; Heckman et al., 2018; Mogstad et al., 2021; Aryal et al., 2022; Mountjoy, 2022). In addition to the empirical convenience, the linear regression relationship may also be derived from a multivariate normal distribution of $\left(Y_{d},X\right)$, or justified by economic models such as the Cobb-Douglas production function which implies that the logarithm of an output is linearly related to the logarithm of inputs, the capital asset pricing model (CAPM) which implies that the expected return of a portfolio is linearly related to the portfolio risk, and the concavely quadratic utility function which implies a linear demand system for differentiated products (Amir et al., 2017).
With a little abuse of notation, we redenote $U_{d}$ to absorb the intercept $\alpha_{d}$ so that
@@
[…]
@@
Noting that Assumption L implicitly requires the potential outcomes to be continuously distributed and supported on the entire real line, we would like to point out that our identification strategy can be adapted to the case of limited valued outcomes as well by specifying a linear latent index. A detailed discussion is left to Appendix D.
@@
[…]
@@
Assumption NL requires the propensity score function to be nonlinear in continuous covariates in a generalized sense. Note that Assumption NL is sufficient but not necessary for nonlinearity of $\pi_{0}$. Appendix C provides other possible forms of sufficient nonlinearity conditions. The combination of Assumptions L and NL will enable identification based on functional forms in a semiparametric version, which can realize the identification of linear coefficients by exploiting nonlinearity elsewhere in the model. Actually, the nonlinear or nonparametric setting of the treatment equation is typically adopted in the applied literature when semiparametric approaches to estimating the MTE are taken, as illustrated in Carneiro et al. (2011) and summarized in Cornelissen et al. (2016). [^1] We thank an anonymous referee for pointing out this observation. The following example provides an economic intuition that underlies the nonlinearity assumption.
Example 4.
Consider the generalized Roy model where individuals choose treatment if the perceived benefit from treatment is greater than the subjective cost of treatment (Eisenhauer et al., 2015). Specifically,
$D=1\left\{Y_{1}-Y_{0}\geq c\left(X\right)+\epsilon\right\},$ (10)
where $c\left(X\right)$ and $\epsilon$ are the observed and unobserved parts of the cost, respectively. Under Assumption L, the treatment equation (10) becomes $D=1\left\{X^{\prime}\left(\beta_{1}-\beta_{0}\right)-c\left(X\right)\geq U\right\}$, where $U=\epsilon-U_{1}+U_{0}$, therefore $\pi\left(x\right)=F_{U\left|X\right.}\left(\left.x^{\prime}\left(\beta_{1}-\beta_{0}\right)-c\left(x\right)\right|x\right)$, which will be a nonlinear function of $x$ if $c\left(x\right)$ is nonlinear or the conditional distribution function of $U$ given $X=x$ is nonlinear. In economic applications, the cost function $c\left(x\right)$ is typically nonlinear in individual and family characteristics. For instance, the subjective cost of higher education would be U-shaped with respect to the family income, because of the relatively high tuition fee for students with disadvantaged backgrounds and high opportunity cost for students with superior backgrounds. On the other hand, the conditional distribution function $F_{U\left|X\right.}\left(\left.u\right|x\right)$ will be nonlinear in $x$ in general if $U$ is not independent of $X$, as shown by Example 6. More generally, the treatment equation (10) is a special case of a utility maximization model of treatment defined as $D=1\left\{\mathfrak{u}_{1}\left(Y_{1},X,\epsilon_{1}\right)\geq\mathfrak{u}_{0}\left(Y_{0},X,\epsilon_{0}\right)\right\}$, where the utility functions $\mathfrak{u}_{1}$ and $\mathfrak{u}_{0}$ are quasi-linear with respect to the first argument. An early version of Lee and Park (2023) gave further examples illustrating why the nonlinearity of the utility functions makes sense in economic studies.
The nonlinearity assumption has two different forms, depending on the number of continuous covariates. When two or more continuous covariates are available, Assumption NL2 will require some variation in $\pi_{0}$ to distinguish it from a single-index function. Concretely, NL2 will not hold if $\pi_{0}\left(x^{C}\right)=f\left(x^{C\prime}\gamma\right)$ for a smooth function $f$, because in this case, both sides of the inequality in (v) are equal to $\left.\gamma_{k}\right/\gamma_{j}$. Otherwise, however, it is difficult to construct examples that violate (v). Therefore, we can informally test NL2 by constructing a test for the single-index specification of the propensity score against a general nonparametric alternative (e.g., Fan and Li, 1996; Stute and Zhu, 2005; Xia, 2009; Maistre and Patilea, 2019). In practice, NL2 can be fulfilled even when $\pi_{0}$ is single-index specified, if interaction or/and quadratic terms are added, as shown by the following Example.
@@
Assumption NL2 is a form of the nonlinearity assumption of Escanciano et al. (2016) that is specific to our model, and thus inherits a demerit that it requires the existence of at least two continuous covariates. However, in empirical studies based on survey data, most of the demographic characteristics are documented as discrete or categorical variables, such as age, gender, race, marital status, educational attainment, and so on. Therefore, we also impose Assumption NL1, as a supplement to NL2, to take into account the situation in which only one continuous covariate is available. NL1 requires $\pi_{0}$ to have at least one stationary point. Namely, NL1 holds if the probability of receiving treatment is unaffected by some local change in the continuous covariate. It’s a differential version of the local irrelevance assumption imposed in nonseparable models for attaining point identification (e.g., Torgovitsky, 2015; D’Haultfœuille et al., 2024). Note that NL1 will hold if $\pi_{0}$ is flat everywhere, i.e., if the continuous covariate has no influence on the treatment choice conditional on the benchmark value of $X^{D}$. However, this extreme case will be ruled out by the subsequent Assumption S. Given that, NL1 implies that $\pi_{0}$ is necessarily nonlinear. NL1 can be extended to the case of no continuous covariate, as discussed in Appendix C. However, using only discrete covariates provides not much identifying variation, which may lead to poor performance in the subsequent model estimation (Garlick and Hyman, 2022). Hence, we focus on the case of at least one continuous covariate. NL1 can be visually tested by plotting the estimated propensity score as a function of the continuous covariate given a benchmark value of $X^{D}$, as illustrated in Figure 1 in the subsequent empirical application.
@@
$D=1\left\{\frac{X^{\prime}\gamma}{\sigma\left(X\right)}\geq\tilde{U}\right\}=1\left\{F_{\tilde{U}}\left(\frac{X^{\prime}\gamma}{\sigma\left(X\right)}\right)\geq V\right\},$
[…]
Unlike in the binary response model, the ubiquitous heteroscedasticity and endogeneity benefit our results while inducing no trouble, because the identification of MTE is irrelevant to the structural coefficients in $\gamma$. Our MTE is defined by the reduced-form treatment error; thus, all we need from the treatment equation is the propensity score, which has a reduced-form nature.
@@
Theorem 1 is an adaption of the identification result of Escanciano et al. (2016). Since Escanciano et al. (2016) considered a general double index model with an unknown link function, their identification result is only up to scale and thus requires a scale normalization assumption. In comparison, Theorem 1 can identify the magnitude of the linear coefficients without any normalization. The proof of Theorem 1 is delegated to Appendix B. Briefly speaking, the proof is grounded on the observed regression functions (8), which summarize the information from the data. As $g_{d}$ is unknown, we need to eliminate it through some subtraction to realize the identification of $\beta_{d}$. When only one continuous covariate exists, the subtraction can be carried out locally around the point satisfying Assumption NL1. Otherwise, our strategy is to perturb two continuous covariates, such as $x_{k}^{C}$ and $x_{j}^{C}$, in such a way that $\pi\left(x\right)$ remains unchanged. Specifically, for each group $d$, we increase $x_{k}^{C}$ by a small $\epsilon$ and simultaneously change $x_{j}^{C}$ by $\epsilon$ multiplied by $\left.-\partial_{k}\pi_{0}\left(x^{C}\right)\right/\partial_{j}\pi_{0}\left(x^{C}\right)$, resulting in a perturbed value of the regression function. Subtracting the perturbed regression function from the original (8) will cancel out $g_{d}$ due to the equality of $\pi\left(x\right)$, giving rise to an equation for $\beta_{d,k}^{C}$ and $\beta_{d,j}^{C}$. Note that the multiplier $\left.-\partial_{k}\pi_{0}\left(x^{C}\right)\right/\partial_{j}\pi_{0}\left(x^{C}\right)$ is the partial derivative of $x_{j}^{C}$ with respect to $x_{k}^{C}$ if we view $x_{j}^{C}$ as an implicit function of the other continuous covariates by equating $\pi_{0}\left(x^{C}\right)$ to a constant. Accordingly, Assumption NL2.(v) implies two linearly independent equations, ensuring an exact solution (i.e., identification) of $\beta_{d,k}^{C}$ and $\beta_{d,j}^{C}$.
It is worth mentioning that our strategy naturally features over-identification in the sense that if there is a pair of points satisfying Assumption NL2, then there must be infinitely many pairs of points for which NL2 holds, because $X^{C}$ is continuously distributed. Moreover, there will generally be more than one value of $X^{D}$ that satisfies Assumptions S, as illustrated by the application in Section 5. Such values of $X^{D}$ can be taken as alternative benchmarks, under which Assumption NL may hold for another set of (pairs of) points of $X^{C}$. Consequently, the identification can be represented as the average of solutions over all (pairs of) points of $X^{C}$ satisfying Assumption NL and over all values of $X^{D}$ satisfying Assumptions S.
@@ 4 Estimation
Our identification strategy for the IV-free MTE implies a separate estimation procedure that works with the partially linear regression (8) by each treatment status. In particular, we recommend the kernel-weighted pairwise difference estimation method proposed by Ahn and Powell (1993) because of its computational simplicity, well-established asymptotic properties, and, most importantly, its capacity to effectively leverage the overidentifying information that underlies the data. Suppose that $\left\{\left(Y_{i},D_{i},X_{i}\right):i=1,2,\cdots,n\right\}$ is a random sample of observations on $\left(Y,D,X\right)$. In the first step, we estimate the nonparametrically specified propensity score using the kernel method, that is,
$\hat{\pi}\left(x\right)=\frac{\sum_{i=1}^{n}D_{i}\left[\prod_{l=1}^{\dim\left(X^{C}\right)}k_{1}\left(\left.\left(X_{il}^{C}-x_{l}^{C}\right)\right/h_{1l}\right)\right]1\left\{X_{i}^{D}=x^{D}\right\}}{\sum_{i=1}^{n}\left[\prod_{l=1}^{\dim\left(X^{C}\right)}k_{1}\left(\left.\left(X_{il}^{C}-x_{l}^{C}\right)\right/h_{1l}\right)\right]1\left\{X_{i}^{D}=x^{D}\right\}},$ (12)
and construct $\hat{P}_{i}=\hat{\pi}\left(X_{i}\right)$ by leaving the $i$-th observation out in the estimation, where $h_{1l}$, $l=1,2,\cdots,\dim\left(X^{C}\right)$, are bandwidths and $k_{1}$ is a univariate kernel function. If the dimension of $X^{D}$ is large, then a smoothed kernel for discrete covariates (Racine and Li, 2004) can be applied as a substitute for the indicator function, to alleviate the potential problem of inadequate observations in each data cell divided by the support of $X^{D}$. If the number of continuous covariates is not small either, then the well-known curse of dimensionality will appear, and a linear-index specification may thus be practically more relevant when modeling the propensity score. The index should include a series of interaction terms and quadratic or even higher-order terms of continuous covariates to supply sufficient nonlinear variation. The linear-index propensity score can be estimated by parametric probit/logit or semiparametric methods (e.g., Powell et al., 1989; Ichimura, 1993; Klein and Spady, 1993; Lewbel, 2000), depending on the distributional assumption on the error term.
@@
$\hat{\omega}_{dij}=1\left\{D_{i}=D_{j}=d\right\}\frac{1}{h_{2}}k_{2}\left(\frac{\hat{P}_{i}-\hat{P}_{j}}{h_{2}}\right),$
@@
$\hat{w}_{di}\left(p\right)=1\left\{D_{i}=d\right\}\frac{1}{h_{3}}k_{3}\left(\frac{\hat{P}_{i}-p}{h_{3}}\right).$
Finally, we plug $\hat{\beta}_{d}$, $\hat{g}_{d}$, $\hat{g}_{d}^{\left(1\right)}$, and $\hat{\pi}$ into identification equations (9) and (11) to estimate the IV-free MTE and other causal parameters, as follows:
$\hat{\Delta}^{\text{MTE}}\left(x,v\right)$ $=$ $x^{\prime}\left(\hat{\beta}_{1}-\hat{\beta}_{0}\right)+\left[\hat{g}_{1}\left(v\right)-\hat{g}_{0}\left(v\right)\right]+v\hat{g}_{1}^{\left(1\right)}\left(v\right)+\left(1-v\right)\hat{g}_{0}^{\left(1\right)}\left(v\right),$ (14)
@@
Theorem 2.
Suppose the assumptions of Theorem 1 hold. Further suppose Assumptions E.1-E.4 given in Appendix E hold. Then for any interior point $v$ of the support of $P$ and any $x$ satisfying $0<\pi\left(x\right)<1$, we have
$\sqrt{nh_{3}^{3}}\left[\hat{\Delta}^{\text{MTE}}\left(x,v\right)-\Delta^{\text{MTE}}\left(x,v\right)\right]$ $\rightarrow$ $N\left(0,\sigma_{\text{MTE}}^{2}\left(v\right)\right),$
$\sqrt{nh_{3}}\left[\hat{\Delta}^{\text{ATE}}\left(x\right)-\Delta^{\text{ATE}}\left(x\right)-b_{\text{ATE}}h_{3}^{2}\right]$ $\rightarrow$ $N\left(0,\sigma_{\text{ATE}}^{2}\right),$
$\sqrt{nh_{3}}\left[\hat{\Delta}^{\text{TT}}\left(x\right)-\Delta^{\text{TT}}\left(x\right)-b_{\text{TT}}\left(\pi\left(x\right)\right)h_{3}^{2}\right]$ $\rightarrow$ $N\left(0,\sigma_{\text{TT}}^{2}\left(\pi\left(x\right)\right)\right),$
$\sqrt{nh_{3}}\left[\hat{\Delta}^{\text{TUT}}\left(x\right)-\Delta^{\text{TUT}}\left(x\right)-b_{\text{TUT}}\left(\pi\left(x\right)\right)h_{3}^{2}\right]$ $\rightarrow$ $N\left(0,\sigma_{\text{TUT}}^{2}\left(\pi\left(x\right)\right)\right),$
$\sqrt{nh_{3}}\left[\hat{\Delta}^{\text{LATE}}\left(x,v_{1},v_{2}\right)-\Delta^{\text{LATE}}\left(x,v_{1},v_{2}\right)-b_{\text{LATE}}\left(v_{1},v_{2}\right)h_{3}^{2}\right]$ $\rightarrow$ $N\left(0,\sigma_{\text{LATE}}^{2}\left(v_{1},v_{2}\right)\right),$
where the asymptotic variances $\sigma_{\text{MTE}}^{2},\cdots,\sigma_{\text{LATE}}^{2}$ are defined in (E.21)-(E.25), respectively, and the asymptotic bias terms are defined as
$b_{\text{ATE}}$ $=$ $\frac{\kappa_{2}}{2}\left(g_{1}^{\left(2\right)}\left(1\right)-g_{0}^{\left(2\right)}\left(0\right)\right),$
$b_{\text{TT}}\left(p\right)$ $=$ $\frac{\kappa_{2}}{2}\left(g_{1}^{\left(2\right)}\left(p\right)+\frac{\left(1-p\right)g_{0}^{\left(2\right)}\left(p\right)-g_{0}^{\left(2\right)}\left(0\right)}{p}\right),$
$b_{\text{TUT}}\left(p\right)$ $=$ $\frac{\kappa_{2}}{2}\left(\frac{g_{1}^{\left(2\right)}\left(1\right)-pg_{1}^{\left(2\right)}\left(p\right)}{1-p}-g_{0}^{\left(2\right)}\left(p\right)\right),$
$b_{\text{LATE}}\left(v_{1},v_{2}\right)$ $=$ $\frac{\kappa_{2}}{2}\left(\frac{v_{2}g_{1}^{\left(2\right)}\left(v_{2}\right)-v_{1}g_{1}^{\left(2\right)}\left(v_{1}\right)+\left(1-v_{2}\right)g_{0}^{\left(2\right)}\left(v_{2}\right)-\left(1-v_{1}\right)g_{0}^{\left(2\right)}\left(v_{1}\right)}{v_{2}-v_{1}}\right),$
with $\kappa_{2}=\int k_{3}\left(u\right)u^{2}du$.
The semiparametric MTE estimator $\hat{\Delta}^{\text{MTE}}\left(x,v\right)$ converges at the same rate $O_{p}\left(\left(nh_{3}^{3}\right)^{-1/2}\right)$ as the local linear estimator for the derivative of a regression function (e.g., Li and Racine, 2007, Theorem 2.7), while the estimators of the aggregated causal parameters converge at the same rate $O_{p}\left(\left(nh_{3}\right)^{-1/2}\right)$ as the kernel estimator for a regression function itself. This is because the estimation of MTE involves plugging in $\hat{g}_{d}^{\left(1\right)}$, the nonparametric derivative estimate, which converges slower than $\hat{g}_{d}$, the nonparametric regression function estimate. Taking the usually applied rule-of-thumb bandwidth $h_{3}\sim n^{-1/5}$ for instance, the MTE estimator will converge at the rate $O_{p}\left(n^{-1/5}\right)$ and the aggregated causal parameters will converge at the rate $O_{p}\left(n^{-2/5}\right)$. Since the slope estimate $\hat{\beta}_{d}$ converges at the parametric rate $O_{p}\left(n^{-1/2}\right)$ that is always faster than $\hat{g}_{d}^{\left(1\right)}$ and $\hat{g}_{d}$, the observed part $x^{\prime}\left(\hat{\beta}_{1}-\hat{\beta}_{0}\right)$ is asymptotically negligible and has no impact on the asymptotic distributions of the estimators. Therefore, for the causal parameters that $x$ does not enter the unobserved part such as the MTE, ATE, and LATE, the asymptotic bias and variance of their estimators will not depend on the observed characteristics.